# A Survey on Multilingual Mental Disorders Detection from Social Media Data

### v1

@@ 6.3 Mental Disorders
Figure 2 shows the distribution of mental disorders in different languages within the datasets. Depression is the most common mental disorder and is well-represented in the data. The languages that lack data on depression are Cantonese, Dutch, Hebrew, Hindi, and Turkish. Suicide is another mental disorder that frequently appears in collections. In contrast, the mental health problems that are least represented include eating disorders, obsessive-compulsive disorder (OCD), attention deficit / hyperactivity disorder (ADHD), autism spectrum disorder (ASD), anxiety, bipolar disorder, and schizophrenia.
@@ 6.4 Annotation Procedure
Most data collections were manually annotated (Figure 2). Manual annotation was carried out by mental health experts or psychologists Narynov et al. (2020); de Oliveira et al. (2022), graduate students who are native speakers of the language of interest Boonyarat et al. (2024); Uddin et al. (2019), or nonexpert individuals. However, some datasets do not specify who the annotators were or what guidelines they followed during the annotation process. Most datasets that collect user-level data from online platforms rely on the self-disclosure of mental health statuses. For example, they rely on explicit mentions of diagnoses (e.g. “I was diagnosed with depression") Tabak and Purver (2020); Villa-Pérez et al. (2023). The third most common annotation method involves asking social media users to complete validated questionnaires to diagnose mental disorders. The most frequently used survey-based methods include the CES-D Tsugawa et al. (2015); Lyu et al. (2023), BDI-II Sun et al. (2022); Stankevich et al. (2019); Ignatiev et al. (2022) or tools specifically designed for certain populations, such as the TMHQ [^10] Thai Mental Health Questionnaire Katchapakirin et al. (2018). Another reliable annotation approach is conducting clinical interviews to assess mental health problems Wołk et al. (2021). Less common and noisier annotation methods include identifying posts based on the presence of specific keywords López-Úbeda et al. (2019), by forum membership Agarwal and Dhingra (2021), or automatic annotation through another model trained on mental health data Cohrdes et al. (2021).
@@ 6.5 Availability of Data Collections
Of the 108 datasets listed in Table 1, only 23 are publicly available for download without any restrictions. These datasets focus on the detection of depression, suicide, and anorexia and are in various languages, including Arabic, Bengali, Brazilian Portuguese, Chinese, Hebrew, Hindi, Spanish, Russian, Roman Urdu, and Thai. For 15 of the datasets, access can be obtained by contacting the authors of the respective research papers, while four datasets require users to complete a data agreement to gain access. Additionally, four datasets are unavailable due to the sensitive nature of the data. For the remaining datasets, the research papers do not provide any information on data availability. Details about the availability of data collections can be found in Appendix A, Table LABEL:tab:datasets.
@@ 7 Mental Disorders Detection Approaches
In this section, we present the NLP methods proposed for the datasets in Section 6. Most approaches are monolingual and specifically target only one non-English language.
Classical approaches
Most approaches use Bag-of-Words, TF-IDF, or Word2Vec for text representation, which are then used as input for classical machine learning models Almouzini et al. (2019); Alghamdi et al. (2020); Helmy et al. (2024) or deep learning models Mann et al. (2020); Tasnim et al. (2022); Ghosh et al. (2023).
Pre-trained transformer-based models
While multilingual models like XLM-Roberta and Multilingual BERT demonstrate strong performance in downstream tasks, only two studies focus exclusively on these models Kabir et al. (2022); Hoque and Salma (2023). In contrast, twelve of the papers in Section 6 rely on pre-trained monolingual models specific to the target language, such as Chinese BERT Yao (2024), AraBERT Abdulsalam et al. (2024), German BERT Zanwar et al. (2023), Bangla BERT Chowdhury et al. (2024) and others. In addition, seven research papers evaluate both language-adapted and multilingual models Hacohen-Kerner et al. (2022); Oliveira et al. (2024).
Translation
Zahran et al. (2025) presented a comprehensive evaluation of LLMs on Arabic data related to depression, suicidal ideation, anxiety, and others. The authors found that LLMs performed better on original Arabic datasets compared to data that had been translated into English. Other works also rely on the detection using data translated from the target language to English Vajrobol et al. (2023). However, Schoene et al. (2025) has shown that automatically translating suicide dictionaries from English to low-resource languages often leads to spelling errors and fails to capture the cultural nuances of the speakers of the target language. When developing mental health models in other languages, some studies rely on translation from English to the target language, such as Greek Skianis et al. (2024) or various Indian languages Rajderkar and Bhat (2024).
Multilingual approaches
@@
Methods developed for multiple languages simultaneously utilize cross-lingual embeddings and make use of information from languages with more mental health-related resources, such as English, to make predictions on Spanish data Coello-Guilarte et al. (2019). Lee et al. (2020) developed a cross-lingual model for suicidal ideation by translating data from Korean to English and Chinese. They used existing dictionaries related to suicidal ideation in these languages to inform predictions on the Korean language.
@@ 8 Cross-cultural and Cross-language Differences in Mental Health Expression
Culture influences the sources of distress, how it is expressed, how it is interpreted, the process of seeking help, and the responses of others Kirmayer et al. (2001). In addition, the way people perceive themselves influences their mental health. In Western cultures, there is a strong emphasis on personal narratives, and people tend to express their emotions more openly, a trend that is reflected in online posts Tokunaga (2009). In contrast, in Asian societies, individuals often internalize their emotional struggles or express them indirectly, influenced by their collectivist values Broczek et al. (2024). Although negative self-thoughts are a common characteristic of depression, in East Asian contexts, self-criticism is often viewed as a sign of healthy functioning Gotlib and Hammen (2008).
@@
Cultural differences in the interpretation of mental health symptoms can lead individuals of certain backgrounds to minimize the psychological effects of mental distress. Instead, they may report more socially acceptable somatic symptoms Kirmayer et al. (2001). Somatic symptoms are common across various cultures, but the ways in which they are reported or understood can differ. In addition, there are culturally specific idioms of distress associated with mental disorders. One such example is the term “nervios” (translated as “nerves” in English), which is a syndrome of distress primarily studied in Latin American communities. This syndrome manifests with psychological and somatic symptoms and has a high comorbidity with anxiety and mood disorders De Snyder et al. (2000). The DSM-V American Psychiatric Association (2013), which is used for the assessment of mental disorders, includes cultural concepts of distress to help clinicians recognize how individuals from various cultures express psychological issues.
@@
The tendency for self-focused attention, often referred to as “I”-language, is considered one of the strongest predictors of depression in language Mihalcea et al. (2024). As a result, the frequency of the pronoun “I” has been used in previous studies as a feature for detecting depression in English. However, it is crucial to carefully consider the applicability of this marker to non-English languages. This association has not been observed in non-Western individuals Rai et al. (2024) or in speakers of Chinese Lyu et al. (2023) or Romanian Trifu et al. (2024). While the pronoun “I” serves as a significant indicator of depression in English, its usage in other languages requires special attention due to linguistic differences. For example, English requires nouns or pronouns to be explicitly included as subjects in sentences. In contrast, some languages, such as Chinese and Romanian, are pro-drop languages, which allow the subject of the action to be omitted Koeneman and Zeijlstra (2019). This can result in a lower frequency of the personal pronoun “I” in these languages.
@@
Indicators of mental disorders are often displayed through metaphors. Depression is often described as weight, pressure, or darkness, and is often portrayed using containment metaphors Charteris-Black (2012). Metaphors are often used by individuals to articulate their experience and psychologists in the therapeutic process Mould et al. (2010). Mental illness metaphors have been extensively studied in English Charteris-Black (2012); Lazard et al. (2016) and have been used to predict mental states Shi et al. (2021); Zhang et al. (2021). With the exception of research in Spanish Coll-Florit and Climent (2023), there is a notable lack of resources to understand metaphors of mental illness in other languages.
It is essential to consider the various cultural and multilingual differences when developing automated methods to predict mental disorders based on language. These differences may explain why many studies have shown that models designed to predict mental illnesses often fail to generalize Aguirre et al. (2021); Abdelkadir et al. (2024).
@@ 9 Research Gaps
As presented in Section 6, most data collection in non-English languages are often from mid- and high-resourced languages, with the exception of Cantonese, Norwegian, and Sinhala. Currently, many languages remain underrepresented, including high-resourced languages like French and mid-to-high resource languages such as Finnish, Croatian, and Vietnamese. Moreover, there is a lack of data collections for low-resource languages, which may hinder the development of online screening tools for individuals who speak these languages. Although few studies have used automatic translation for building datasets in languages other than English, it cannot accurately capture the cultural nuances of native speakers of the target language Schoene et al. (2025).
Cross-lingual expressions in underrepresented mental disorders
Although there are mental health-related datasets available in non-English data, most of them primarily focus on depression and suicide. Other mental disorders, such as anxiety, OCD, bipolar disorder, and PTSD, are underrepresented. To gain a better understanding of how these disorders manifest in the online language, the research community needs more linguistically diverse collections that encompass a wider range of mental disorders. This approach would not only facilitate a broader exploration of mental health expressions in various languages, but also help develop more inclusive and effective online mental health screening tools worldwide.
@@
As highlighted in Section 7, most NLP approaches have focused on processing data in a single target language, with multilingual approaches addressing multiple languages being almost nonexistent. Most existing NLP models developed for mental disorders detection do not support multiple languages effectively, which limits their applicability in multicultural and multilingual settings where mental health issues may manifest differently.
@@
Although most of the datasets presented in this paper rely on manual annotation for labeling the data related to mental disorders, it is often unclear who did the annotations. The authors of the research papers should provide specific details about the annotation process, such as whether the annotators are mental health experts or non-experts, if they are native speakers of the target language, and whether they understand the cultural differences in the manifestations of mental disorders. These factors significantly impact the quality and reliability of the data, as understanding cultural nuances is essential in interpreting mental health expressions.
@@
While many mental health studies in English emphasize the importance of explainable approaches Yang et al. (2023a); Souto et al. (2023); Yang et al. (2023b), there is a significant opportunity for applying explainable approaches to non-English languages. Currently, few studies have examined model explainability in Bengali Ghosh et al. (2023) and Thai Vajrobol et al. (2023). These methods may help in understanding the various manifestations of mental disorders.
@@ 10 Conclusion
In this paper, we presented a comprehensive review of research for mental disorders detection from multilingual data sourced from social media. We highlight cross-cultural and multilingual differences in mental health expressions and provide a comprehensive list of data collections that can be used to develop multilingual NLP models for online mental health screening. Our focus was on non-English resources, as most previous research has focused on English Skaik and Inkpen (2020); Harrigian et al. (2021). Lastly, we identified several gaps in current research that we hope will be addressed in future interdisciplinary studies.
Future Directions and Call to Action
We aim to encourage researchers to develop mental health datasets in low-resource languages, fostering interdisciplinary collaborations with experts from psychology and mental health organizations, as seen in successful previous projects like REMO COST Action [^11] https://projects.tib.eu/remo, and PsyMine Ellendorff et al. (2016), which have primarily focused on English. By involving community members, multilingual shared tasks can be organized to identify mental disorders across different languages, inspired by successful SemEval multilingual tasks for offensive language Zampieri et al. (2020) and emotion detection Muhammad et al. (2025). Researchers can work together to annotate data in underrepresented languages while adhering to ethical protocols. By participating in these tasks, members of the ACL community can gain access to data collections that are essential for developing multilingual models. Such initiatives will improve the visibility of multilingual mental disorder detection and encourage further collaborations, providing researchers with more opportunities to address challenges in this field. Researchers can focus on building data collections for underrepresented mental disorders beyond depression and suicide, adhering to ethical guidelines and providing transparency in the annotation process Benton et al. (2017). Recent advances in explainability can also be applied to better understand the cultural manifestations of mental disorders.
@@
Our paper aims to provide a comprehensive review of cross-cultural language differences and the datasets available for developing multilingual NLP models. We included 108 data collections in this study and carefully reviewed each paper cited in our survey. However, it is possible that we may have overlooked some works that do not explicitly mention in their title or abstract that they focus on non-English languages.
@@
There are ongoing concerns regarding the construct validity of models trained on data collected from social media, specifically whether these models effectively measure the manifestations of mental disorders Chancellor and De Choudhury (2020). The datasets used in this survey predominantly rely on manual annotation or labeling through validated questionnaires, which are considered more reliable methods for annotation. However, it is essential to conduct interdisciplinary research and ground the constructs being measured in both theoretical and clinical frameworks. For example, clinical depression (or major depressive disorder) is fundamentally different from merely “feeling depressed.” The latter may refer to temporary feelings, while clinical depression encompasses a range of persistent symptoms. These symptoms may include depressed mood, loss of interest in previously enjoyed activities, changes in body weight, sleep disturbances, fatigue, psychomotor agitation or retardation, feelings of guilt, and thoughts of death or suicidal ideation American Psychiatric Association (2013). To be diagnosed with depression, these symptoms must be persistent and significantly impair an individual’s ability to function. Prioritizing interdisciplinary collaboration and rigorous validation methods is essential in addressing the complexities of mental health.

### v2

@@ 5.3 Mental Disorders
Figure 3 shows the distribution of mental disorders in different languages within the datasets. Depression is the most common mental disorder and is well-represented in the data. The languages that lack data on depression are Cantonese, Dutch, Hebrew, Hindi, and Turkish. Suicide is another mental disorder that frequently appears in collections. In contrast, the mental health problems that are least represented include eating disorders, obsessive-compulsive disorder (OCD), attention deficit / hyperactivity disorder (ADHD), autism spectrum disorder (ASD), anxiety, bipolar disorder, and schizophrenia. Hazell et al. (2022)
@@ 5.4 Annotation Procedure
Most data collections were manually annotated (Figure 3). Manual annotation was carried out by mental health experts or psychologists Narynov et al. (2020); de Oliveira et al. (2022), graduate students who are native speakers of the language of interest Boonyarat et al. (2024); Uddin et al. (2019), or nonexpert individuals. However, some datasets do not specify who the annotators were or what guidelines they followed during the annotation process. Most datasets that collect user-level data from online platforms rely on the self-disclosure of mental health statuses. For example, they rely on explicit mentions of diagnoses (e.g. “I was diagnosed with depression") Tabak and Purver (2020); Villa-Pérez et al. (2023). The third most common annotation method involves asking social media users to complete validated questionnaires to diagnose mental disorders. The most frequently used survey-based methods include the CES-D Tsugawa et al. (2015); Lyu et al. (2023), BDI-II Sun et al. (2022); Stankevich et al. (2019); Ignatiev et al. (2022) or tools specifically designed for certain populations, such as the Thai Mental Health Questionnaire (TMHQ) Katchapakirin et al. (2018). Another reliable annotation approach is conducting clinical interviews to assess mental health problems Wołk et al. (2021). Less common and noisier annotation methods include identifying posts based on the presence of specific keywords López-Úbeda et al. (2019), by forum membership Agarwal and Dhingra (2021), or automatic annotation through another model trained on mental health data Cohrdes et al. (2021).
Manual annotation and annotation based on self-disclosure of diagnosis rely on individuals’ willingness to disclose personal information on social media. Research shows that collectivist cultures tend to have lower rates of self-disclosure compared to individualist cultures in online contexts Tokunaga (2009). However, there are notable differences within these cultures. For example, although Koreans are less likely to share personal information, they often provide more in-depth disclosures compared to individuals in the United States Yoo (2012). In addition, self-disclosure annotation relies on individuals explicitly naming their conditions, which may not occur in cultures where mental health is stigmatized. These annotation choices can have significant consequences for model behavior, potentially resulting in systematic false negatives for populations whose expressions of distress do not align with dominant annotation frameworks.
@@ 5.5 Availability of Data Collections
Out of the 108 datasets listed in Table 1, only 23 are publicly available for download without restrictions. This is not surprising, as mental health datasets are often subject to privacy, ethical, and legal concerns, unlike many other NLP domains. These open datasets focus on the detection of depression, suicide, or anorexia in various languages, including Arabic, Bengali, Brazilian Portuguese, Chinese, Hebrew, Hindi, Spanish, Russian, Roman Urdu, and Thai. For 15 of the datasets, access can be obtained by contacting the authors of the respective research papers, while four datasets require users to complete a data agreement to gain access. In addition, four datasets are unavailable due to the sensitive nature of the data. For the remaining datasets, the research papers do not provide any information on data availability. Details about the availability of data collections can be found in Appendix A, Table LABEL:tab:datasets.
@@ 6 Mental Disorders Detection Approaches
In this section, we present the methods proposed for the datasets in Section 5. Most approaches are monolingual and specifically target only one non-English language. In Appendix A, Table LABEL:tab:datasets, we include detailed information on the methods used and evaluation scores for all the datasets.
Classical Monolingual Approaches
Most approaches rely on feature-based text representations such as Bag-of-Words, TF-IDF, and Word2Vec, combined with traditional classifiers like SVM, Logistic Regression Almouzini et al. (2019); Alghamdi et al. (2020); Helmy et al. (2024) or deep learning models Mann et al. (2020); Tasnim et al. (2022); Ghosh et al. (2023). These methods are straightforward and can be adapted to any language, making them suitable for the detection task.
Monolingual pre-trained models
With the rise of large pre-trained transformers, language-specific models such as Chinese BERT Yao (2024), AraBERT Abdulsalam et al. (2024), German BERT Zanwar et al. (2023), Bangla BERT Chowdhury et al. (2024), BERTimbau Santos et al. (2024), IndicBERT Agarwal and Dhingra (2021) are being used for detecting mental disorders. However, a significant limitation of these pre-trained transformer-based models is that they are available only for a limited number of high- and mid-resource languages, rendering them impractical for use with low-resource languages.
Multilingual models
@@
Methods developed for multiple languages simultaneously are rare. Some approaches use cross-lingual embeddings and leverage information from languages with more extensive mental health-related resources, such as English, to make predictions on Spanish data Coello-Guilarte et al. (2019). Lee et al. (2020) developed a cross-lingual model for suicidal ideation by translating data from Korean to English and Chinese. Although multilingual pre-trained models like XLM-Roberta and Multilingual BERT are frequently used, they often perform worse than language-specific models, such as BERTimbau and HeBERT Oliveira et al. (2024); Hacohen-Kerner et al. (2022). This discrepancy suggests that adapting models to specific cultural and linguistic contexts is often more effective than relying solely on general multilingual pre-training. Consequently, relying on a single model for multiple languages can increase the likelihood of errors.
Translation-based
Zahran et al. (2025) conducted a comprehensive evaluation of LLMs on Arabic data related to depression, suicidal ideation, anxiety, and other mental disorders. The authors showed that LLMs performed better on original Arabic datasets than on data that had been translated into English. Some studies also use data translated from the target language to English Vajrobol et al. (2023). When developing mental health models in languages other than English, certain studies use translations from English to the target language, as seen with Greek Skianis et al. (2024) and various Indian languages Rajderkar and Bhat (2024). However, Schoene et al. (2025) has shown that automatically translating suicide-related dictionaries from English to low-resource languages often results in spelling errors and fails to capture the cultural nuances of speakers in the target language.
LLMs
While LLMs have been successfully used in mental health tasks in English Wang et al. (2024b); Shah et al. (2025), research for other languages is limited. For instance, Zahran et al. (2025) evaluated LLMs on Arabic data and found that slight changes in instructions could harm performance, leading to parsing errors and prompt adherence issues. There are also concerns about increased hallucinations Lauscher et al. (2025) and unsafe responses Wang et al. (2024b) in languages other than English. Consequently, the use of LLMs in mental health tasks for non-English languages should be approached with caution.
Performance across different languages
As mentioned in Section 3, the detection of mental disorders is primarily approached as a classification task and evaluated using standard metrics such as Accuracy, Precision, Recall, and F1-score. Detailed performance scores can be found in Table LABEL:tab:datasets in Appendix A. The availability of resources has a significant impact on the performance of supervised or pre-trained language models. High-resource languages, such as Japanese, Chinese, and Arabic, generally yield strong results, indicating that larger and well-annotated datasets contribute to more reliable detection. In contrast, some languages, such as Sinhala, Brazilian Portuguese, and Bengali, show highly variable performance when comparing the detection of depression and suicidal ideation. For other languages included in this survey, such as French, German, Russian, and Italian, the performance in detecting depression is relatively low. Current methods demonstrate unequal effectiveness, which may leave speakers of certain languages without appropriate mental health screening tools. The success of these models is influenced by factors such as language-specific methodologies, the quality of datasets, the availability of resources, and cultural differences in expression. In addition, performance may vary depending on whether the task is binary classification, multi-class classification, or regression. Performance can also vary depending on whether the prediction is based on a single post (post-level classification) or multiple posts from the same user (user-level classification).
@@ 7 Cross-cultural and Cross-language Differences in Mental Health Expression
Culture influences the sources of distress, how it is expressed, how it is interpreted, the process of seeking help, and the responses of others Kirmayer et al. (2001). In addition, the way people perceive themselves influences their mental health. In Western cultures, there is a strong emphasis on personal narratives, and people tend to express their emotions more openly, a trend that is reflected in online posts Tokunaga (2009). In contrast, in Asian societies, individuals often internalize their emotional struggles or express them indirectly, influenced by their collectivist values Hu et al. (2014). Although negative self-thoughts are a common characteristic of depression, in East Asian contexts, self-criticism is often viewed as a sign of healthy functioning Gotlib and Hammen (2008).
@@
Cultural differences in the interpretation of mental health symptoms can lead some individuals to minimize psychological distress, often reporting socially acceptable somatic symptoms such as body pain or fatigue Kirmayer et al. (2001). While somatic symptoms are common across cultures, their expression and understanding can vary. Culturally specific idioms of distress, such as ”nervios” [^11] Translated as “nerves” in English. in Latin American communities, illustrate this, as they encompass both psychological and somatic symptoms and show a high comorbidity with anxiety and mood disorders De Snyder et al. (2000). The DSM-V APA (2013) includes cultural concepts of distress to aid clinicians in recognizing diverse expressions of psychological issues. In addition, Rai et al. (2025) found that Indian social media users on Reddit are more focused on seeking help than on discussing their symptoms or diagnoses.
@@
The tendency for self-focused attention, often referred to as “I”-language, is considered one of the strongest predictors of depression in language Mihalcea et al. (2024). As a result, the frequency of the pronoun “I” has been used in previous studies as a feature for detecting depression in English. However, it is crucial to carefully consider the applicability of this marker to non-English languages. This association has not been observed in non-Western individuals Rai et al. (2024) or in speakers of Chinese Lyu et al. (2023) or Romanian Trifu et al. (2024). While the pronoun “I” serves as an indicator of depression in English, its usage in other languages requires special attention due to linguistic differences. For example, English requires nouns or pronouns to be explicitly included as subjects in sentences. In contrast, some languages, such as Chinese and Romanian, are pro-drop languages, which allow the subject of the action to be omitted Koeneman and Zeijlstra (2019), resulting in a lower frequency of the personal pronoun “I” in these languages.
@@
Indicators of mental disorders are often displayed through metaphors. Depression is often described as weight, pressure, or darkness, and is often portrayed using containment metaphors Charteris-Black (2012). Metaphors are used by individuals to articulate their experiences, and psychologists use them in the therapeutic process Mould et al. (2010). Mental illness metaphors have been extensively studied in English Charteris-Black (2012); Lazard et al. (2016) and have been used to predict mental states Shi et al. (2021); Zhang et al. (2021). With the exception of research in Spanish Coll-Florit and Climent (2023), there is a lack of resources to understand metaphors of mental illness beyond English.
Considering cultural and multilingual differences is crucial when developing automated methods for predicting mental disorders based on language. These variations may explain the failure of many predictive models to generalize effectively Aguirre et al. (2021); Abdelkadir et al. (2024).
@@ 8 Research Gaps
Linguistic features used in English do not transfer reliably across languages
Many commonly used linguistic markers of mental disorders, such as the use of first-person singular pronouns or sentiment polarity, are often considered universal features. However, these markers are actually dependent on language and shaped by cultural contexts, as previous research has indicated Trifan et al. (2020); Rai et al. (2024). While many studies have investigated markers of depression in language across cultures, most have focused primarily on individuals from various cultures who communicate in English Aguirre et al. (2021); Abdelkadir et al. (2024); Rai et al. (2025). There is a significant lack of research on how cross-cultural factors influence the expression of mental disorders in individuals’ native languages on social media, as well as how these factors affect the identification of such disorders.
@@
As presented in Section 5, most datasets are often from mid- and high-resourced languages, with the exception of Cantonese, Norwegian, and Sinhala. Currently, many languages remain underrepresented, including high-resource languages like French and mid-to-high-resource languages such as Finnish, Croatian, and Vietnamese. Moreover, there is a lack of datasets in low-resource languages, which may hinder the development of online screening tools for individuals who speak these languages. While some studies have attempted to use automatic translation to build datasets in languages other than English, this approach often fails to accurately capture the cultural nuances of native speakers of the target language Schoene et al. (2025).
@@
As highlighted in Section 6, most NLP approaches have focused on processing data in a single target language, with multilingual approaches being almost nonexistent. Most existing NLP models developed for mental disorders detection do not support multiple languages effectively, which limits their applicability in multicultural and multilingual settings where mental health issues may manifest differently. While multilingual and cross-lingual approaches are often proposed as solutions to data scarcity, generalizability failures may stem not only from resource limitations but also from unmodeled cultural variation.
@@
Although most of the datasets presented in this paper rely on manual annotation for labeling the data related to mental disorders, it is often unclear who performed the annotations. The authors of the research papers should provide specific details about the annotation process, such as whether the annotators are mental health experts or non-experts, if they are native speakers of the target language, and whether they understand the cultural differences in the manifestations of mental disorders. These factors significantly impact the quality and reliability of the data, as understanding cultural nuances is essential in interpreting mental health expressions. For annotations based on questionnaires, it is crucial to use validated questionnaires specific to the target language rather than general English versions. Previous research has demonstrated that incorporating information about cultural idioms of distress in psychological assessments for mental disorders can enhance their validity Cork et al. (2019). Therefore, such considerations should also be integrated into the annotation process for mental disorders across cultures.
@@
While many mental health studies in English emphasize the importance of explainable approaches Yang et al. (2023); Souto et al. (2023); Yang et al. (2024), there is a significant opportunity for applying explainable approaches to non-English languages. This could enhance our understanding of how mental disorders manifest culturally in social media language and provide insights into the diverse expressions of mental disorders. Currently, few studies have examined model explainability in Bengali Ghosh et al. (2023) and Thai Vajrobol et al. (2023).
@@ 9 Conclusion
In this paper, we presented a comprehensive survey of research for mental disorders detection in social media data beyond English. We highlight cross-cultural and multilingual differences in mental health expressions and provide a comprehensive list of datasets that can be used to develop multilingual NLP models for online mental health screening. Our focus was on non-English resources, as most previous research has focused on English Skaik and Inkpen (2020); Harrigian et al. (2021). Lastly, we identified several gaps in current research that we hope will be addressed in future interdisciplinary studies.
Future Directions and Calls to Action
We aim to encourage researchers to develop mental health datasets in languages other than English, fostering interdisciplinary collaborations with experts from psychology and mental health organizations, as seen in successful previous projects like REMO COST Action [^12] https://projects.tib.eu/remo, and PsyMine Ellendorff et al. (2016), which have primarily focused on English. By involving community members, multilingual shared tasks can be organized to identify mental disorders across different languages, inspired by successful SemEval multilingual tasks for offensive language Zampieri et al. (2020) and emotion detection Muhammad et al. (2025a, b). Researchers can work together to annotate data in underrepresented languages while adhering to ethical protocols. By participating in these tasks, members of the ACL community can gain access to datasets that are essential for developing multilingual models. Such initiatives will improve the visibility of multilingual mental disorder detection and encourage further collaborations, providing researchers with more opportunities to address challenges in this field. Researchers can focus on building datasets for underrepresented mental disorders beyond depression, adhering to ethical guidelines and providing transparency in the annotation process Benton et al. (2017). Recent advances in explainability can also be applied to better understand the cultural manifestations of mental disorders.
@@
Our paper aims to provide a comprehensive survey of cross-cultural language differences and the datasets available for developing multilingual NLP models. We included 108 data collections in this study and carefully reviewed each paper cited in our survey. However, it is possible that we may have overlooked some works that do not explicitly mention in their title or abstract that they focus on languages other than English.
@@
There are ongoing concerns regarding the construct validity of models trained on data collected from social media, specifically whether these models effectively measure the manifestations of mental disorders Chancellor and De Choudhury (2020). The datasets used in this survey predominantly rely on manual annotation or labeling through validated questionnaires, which are considered more reliable methods for annotation. However, it is essential to conduct interdisciplinary research and ground the constructs being measured in both theoretical and clinical frameworks. For example, clinical depression (or major depressive disorder) is fundamentally different from merely “feeling depressed.” The latter may refer to temporary feelings, while clinical depression encompasses a range of persistent symptoms. These symptoms may include depressed mood, loss of interest in previously enjoyed activities, changes in body weight, sleep disturbances, fatigue, psychomotor agitation or retardation, feelings of guilt, and thoughts of death or suicidal ideation APA (2013). To be diagnosed with depression, these symptoms must be persistent and significantly impair an individual’s ability to function. Prioritizing interdisciplinary collaboration and rigorous validation methods is essential in addressing the complexities of mental health.
@@
Acknowledgments
We sincerely appreciate the insightful feedback provided by the anonymous reviewers, which has greatly contributed to the improvement of this paper.